-3x+4y-5=-14 rewrite the equation so that y is a function of x
Answer by checkley71(8403)

-3X+4Y-5=-14
-3X=-4Y+5-14
-3X=-4Y-9
3X=4Y-9
X=4Y/3-9/3
X=4Y/3-3